A garden hose pointed at an angle of 25 drgrees above the horizontal splashes water on a sunbather lying on the ground 4.4m away in the horizontal direction. If the hose is held 1.4m above the ground at what speed does the water leave the nozzel
1
Expert's answer
2020-09-24T11:07:21-0400


Write the expression for range:


"R=v\text{ cos}\theta\cdot t."

The time is


"t=t_{up}+t_{down}.\\\space\\
t_{up}=\frac{v\text{ sin}\theta}{g},\\\space\\
t_{down}=\sqrt{\frac{2H}{g}}."

The height the water will fall down is


"H=\frac{gt_{up}^2}{2}+h=\frac{v^2\text{sin}^2\theta}{2g}+h."

Thus, the time down becomes


"t_{down}=\sqrt{\frac{2\big(\frac{v^2\text{sin}^2\theta}{2g}+h\big)}{g}}."

Now, express the time from the range (see the first equation):


"t=\frac{R}{v\text{ cos}\theta}=t_{up}+t_{down},\\\space\\
\frac{R}{v\text{ cos}\theta}=\frac{v\text{ sin}\theta}{g}+\sqrt{\frac{2\big(\frac{v^2\text{sin}^2\theta}{2g}+h\big)}{g}}."

All variables are given except the initial velocity. So, we can easily solve this equation:

"v=5.8\text{ m/s}."